# BEAM: Bridging Physically-based Rendering and Gaussian Modeling for Relightable Volumetric Video

### v1

@@ 5 Conclusion
Although our method achieves high-quality immersive rendering, there are some limitations. First, the integration of new attributes into the original Gaussian expression will lead to substantial storage requirements for dynamic Gaussian sequences. Future work will consider compressing these sequences. Second, we approximate the rendering equation to obtain the 2D material maps, which introduces errors in our decoupled material results and does not accurately reflect the real physical world. Future work may address this issue by incorporating large models like video generation. Furthermore, our approach, focusing on relighting for a dynamic reconstruction sequence, does not support pose-driven animation or the generation of new poses. Future work will focus on optimizing these aspects to improve robustness and applicability.
@@
We have presented a Gaussian-based approach for reconstructing detailed geometry and PBR materials to produce relightable volumetric videos. We employ a coarse-to-fine training strategy and effective geometric constraints to accurately model the dynamic geometry of 4D Gaussians. Additionally, we decouple PBR materials by using ray tracing to compute the lighting effects and obtain base color and AO maps, while leveraging generative methods to infer roughness. These materials are then baked into the corresponding attributes of the Gaussians. With deferred shading and ray tracing techniques, our Gaussian sequence supports both efficient real-time rendering and more realistic offline rendering. Experimental results demonstrate the advantages of our approach in generating high-quality dynamic normal maps and material decomposition, as well as its relightability under a variety of lighting conditions. Our method is highly compatible with traditional CG engines, offering significant potential for enhancing rendering realism and flexibility.
Appendix A Energy Term
To model human motion using Gaussian sequences, we first employ the color energy $E_{\text{color}}$ during the training process:
$E_{\text{color}}=(1-\lambda_{\text{color}})\mathcal{L}_{1}+\lambda_{\text{color}}\mathcal{L}_{\text{D-SSIM}},$ (8)
where $\mathcal{L}_{1}$ is the $L_{1}$ photometric loss and $\mathcal{L}_{\text{D-SSIM}}$ is the D-SSIM term. We then employ a smooth regularizer to constrain the joint Gaussians motion locally as-rigidas-possible(ARAP):
$E_{\text{smooth}}=$ $\sum_{i}\sum_{k\in\mathcal{N}(i)}w_{i,k}\left\|R\left(q_{i,t}^{j}*{q_{i,t-1}^{j}}^{-1}\right)\right.$ (9)
$\quad\left.\left(p_{k,t-1}^{j}-p_{i,t-1}^{j}\right)-\left(p_{k,t}^{j}-p_{i,t}^{j}\right)\right\|_{2}^{2},$
where $\mathcal{N}(i)$represents the set of neighboring joint Gaussian kernels of $i$, $R(\cdot)$ converts quaternion back into a rotation matrix and $w_{i,k}$ corresponds to the blending weights according to DualGS [30]. Furthermore, we apply the temporal regularizer $E_{\text{temp}}$ to promote coherent appearances by constraining Gaussian attributes ($C_{i,t},\sigma_{i,t},s_{i,t}$) to be consistent with the previous frame:
$E_{\text{temp}}=\sum_{a\in\{\mathcal{C},\sigma,s\}}\lambda_{a}\|a_{i,t}-a_{i,t-1}\|_{2}^{2},$ (10)
Together with $E_{\text{normal}}$ we mentioned above, the overall energy term in our 4D Gaussian modeling is expressed as:
$E=\lambda_{\text{color}}E_{\text{color}}+\lambda_{\text{smooth}}E_{\text{smooth}}+\lambda_{\text{temp}}E_{\text{temp}}+\lambda_{\text{normal}}E_{\text{normal}},$ (11)
which helps us obtain accurate geometric information while modeling human performance sequences.
Appendix B Base Color Decomposition
To compute base color, we ignore the effect of indirect illumination caused by reflections on the surface of the human body, the rendering equation can be approximated as:
$L_{o}\left(x,\boldsymbol{\omega}_{o}\right)\approx\int_{\Omega}f_{r}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (12)
where $x$ and $\mathbf{n}$ are the surface point and its normal vector, $f_{r}$ is the Bidirectional Reflectance Distribution Function (BRDF), $L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ denotes the outgoing radiance in direction $\boldsymbol{\omega}_{o}$.
We decompose the BRDF term $f_{r}$ into two components, the diffuse term $f_{r_{d}}$ and the specular term $f_{r_{s}}$, where we use the Lambertian model for $f_{r_{d}}$ and Cook-Torrance microfacet specular shading model [8] for $f_{r_{s}}$:
$f_{r}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)$ $=f_{r_{d}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)+f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right),$ (13)
$f_{r_{d}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)$ $=\frac{1-F}{\pi}\rho(x),$ (14)
$f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)$ $=\frac{DGF}{4(\mathbf{n}\cdot\boldsymbol{\omega}_{i})(\mathbf{n}\cdot\boldsymbol{\omega}_{o})},$ (15)
where $D$, $G$, $F$ are respectively the GGX normal distribution function (NDF), the geometric attenuation function, and the approximated Fresnel term. We then split $L_{o}$ into two parts:
$L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\int_{\Omega}\left(f_{r_{d}}+f_{r_{s}}\right)\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i},$ (16)
$=\int_{\Omega}f_{r_{d}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}+\int_{\Omega}f_{r_{s}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i},$ (17)
$=\frac{\rho(x)}{\pi}\int_{\Omega}\left(1-F\right)\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}+\int_{\Omega}f_{r_{s}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i},$ (18)
where $\mathcal{L}=V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})$.
The $D$ and $G$ in $f_{r_{s}}$ are both irrelevant with the base color $\rho$. For the Fresnel term $F$, we use the Schlick Fresnel approximation [52]:
$F=F_{0}+(1-F_{0})\left(1-\left(\mathbf{h}\cdot\boldsymbol{\omega}_{o}\right)\right)^{5},$ (19)
where $\mathbf{h}=\frac{\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}}{\left\|\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}\right\|}$, $F_{0}$ represents the specular reflectance at normal incidence and is achromatic for dielectrics and chromatic for metals. Since the human body is mostly composed of dielectric materials, we assume $F_{0}=0.04$, so that $F$ is also irrelevant with $\rho$. Consequently, we can solve the base color $\rho$ from Eq. 18:
$\rho(x)=\frac{\pi\left(L_{o}(x,\boldsymbol{\omega}_{o})-\int_{\Omega}f_{r_{s}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}\right)}{\int_{\Omega}\left(1-F\right)\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}}.$ (20)
Appendix C Additional Evaluations
In this section, we present more details on the qualitative comparison conducted on synthetic human data, including normal, roughness, base color, AO, and relight rendering effects in Fig. 9.
Appendix D Applications
In Fig. 10, we showcase the application of watching an actor dancing in a scene with varying lighting. Viewers can immerse themselves in a realistic environment with dramatic lighting changes, observing the actor’s dance performance seamlessly integrated into the surroundings. We can capture, process, and generate high-fidelity human 4D assets with our dome cameras and BEAM pipeline, reproducing outstanding performances in different scenes and lightings.
At the same time, our compatibility with the traditional CG pipeline allows us to edit scenes in the Unity plugin we developed or integrate multiple human performances, as shown in Fig. 11. We can align multiple performances through CG platform, edit the scene and lighting, and present stunning visual effects of the performances.
Figure 9: We present additional comparison results on synthetic humans, demonstrating the high-quality relighting capabilities of our method.
Figure 10: We demonstrate our VR application on PICO, enabling an immersive experience with relightable volumetric video.
Figure 11: We showcase our custom Unity plugin for relightable 4D Gaussian assets. Our relightable 4D Gaussians are CG-friendly, supporting scene and lighting environment editing within Unity.

### v2

@@ 6. Conclusion
Although our method achieves high-quality immersive rendering, there are some limitations. First, we approximate the rendering equation to obtain the 2D material maps, which introduces errors in our decoupled material results and does not accurately reflect the real physical world. Future work may address this issue by incorporating large models like video generation. Furthermore, our approach, focusing on relighting for a dynamic reconstruction sequence, does not support pose-driven animation or the generation of new poses. Future work will focus on optimizing these aspects to improve robustness and applicability.
@@
We have presented a Gaussian-based approach for reconstructing detailed geometry and PBR materials to produce relightable volumetric videos. We employ a coarse-to-fine training strategy and effective geometric constraints to accurately model the dynamic geometry of 4D Gaussians. Additionally, we decouple PBR materials by using ray tracing to compute the lighting effects and obtain base color and AO maps, while leveraging generative methods to infer roughness. These materials are then baked into the corresponding attributes of the Gaussians. With deferred shading and ray tracing techniques, our Gaussian sequence supports both efficient real-time rendering and more realistic offline rendering. Experimental results demonstrate the advantages of our approach in generating high-quality dynamic normal maps and material decomposition, as well as its relightability under a variety of lighting conditions. Our method is highly compatible with traditional CG engines, offering significant potential for enhancing rendering realism and flexibility, enabling users to immerse themselves in and interact with dynamic, relightable volumetric worlds.
Acknowledgements.
This work was supported by the National Key R&\&D Program of China (2022YFF0902301), Shanghai Local College Capacity Building Program (22010502800), and Shanghai Frontiers Science Center of Human-centered AI (ShangHAI).